OpenStudy (anonymous):

The total revenue received for the sale of x items is given by R(x) =25 ln (4x+3), while the total cost to produce x items is c(x)=x/3. Find the approx # of items that should be manufactured so profit, r(x)-c(x) is a max. (round answer to nearest integer)

OpenStudy (anonymous):

just my guess but wouldn't it be where rx and cx intersect?

OpenStudy (anonymous):

or would it be profit =rx-cx and find the max of profit?

OpenStudy (anonymous):

I'm not to sure...Would this help? How many items should be manufactured to maximize the profit?

OpenStudy (campbell_st):

create a profit function P(x) = R(x) - C(x) \[P(x) = 25 \ln(4x + 3) - \frac{x}{3}\] differentiate \[P'(x) = \frac{ 100}{4x + 3} - \frac{1}{3}\] let P'(x) = 0 \[\frac{100}{4x + 3} - \frac{ 1}{3} = 0\] \[\frac{100}{4x + 3} = \frac{1}{3} ....... 300 = 4x + 3\] solve for x.